# Self-organizing Multiagent Target Enclosing under Limited Information and Safety Guarantees

### v1

@@ III-B Design of Potential Function
As indicated by the behavior of pursuers observed in Section III, each pursuer must exhibit two distinct behaviors. When there is no nearby colliding pursuer, the $i$th pursuer’s objective is to reach the desired proximity from the target to enclose it. However, if there is a nearest colliding pursuer, the $i$th pursuer needs to maneuver along a loiter circle with a radius larger than the one corresponding to the nearest colliding pursuer to prevent any potential collisions. These factors enable us to break down the complex multiagent target enclosing problem into more manageable two-body (one pursuer and the target) or three-body (two pursuers and the target) scenarios at any given point in time. Consequently, in this subsection, we present the design of a potential function that will serve as the foundation for developing the target enclosing guidance law.
Let’s define the range error for the $i$th pursuer as,
${e}_{i}=r_{iT}-r_{d},$ (22)
where $r_{d}$ is the desired fixed proximity from the target such that $\dot{r}_{d}=0$. This error variable plays a crucial role in the design of the potential function to ensure inter-pursuer safety. To analyze the error dynamics, we differentiate $e_{i}$ with respect to time and use (1a) to obtain the range error rate for the $i$th pursuer as,
$\dot{e}_{i}=\dot{r}_{iT}=-v\cos\sigma_{i}$ (23)
since $\dot{r}_{d}=0$. From the above equation, it is evident that $\dot{e}_{i}<0,$ if $\sigma_{i}\in[0,\pi/2)$, which indicates the $i$th pursuer’s motion toward the target, whereas $\dot{e}_{i}>0,$ represents the opposite case because $\sigma_{i}\in(\pi/2,\pi]$. The $i$th pursuer moves on a circle of fixed radius from the target when $\dot{e}_{i}=0$ when $\sigma_{i}=\pi/2$. Therefore, nullifying the range error given in (22), will ensure that the $i$th pursuer converges to the desired proximity around the target, but additional safety considerations are needed to prevent the inter-pursuer collision.
@@
$U_{i}(e_{i},e_{j})=\lambda_{i}\frac{e_{i}^{2}}{2}+\eta_{i}\delta_{i}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}$ (24)
where $\lambda_{i}$, $\eta_{i}$, and $\Delta_{i}$ are positive scaling constants in the potential function. The subscript $j$ corresponds to $P_{j}\in\mathcal{C}_{i}$, which is the nearest colliding pursuer corresponding to the $i$th pursuer. The variable $\delta_{i}$ is a conditional one such that
$\delta_{i}=\begin{cases}1;\,&\text{if}\,\mathcal{C}_{i}\neq\emptyset~{}(\mathcal{C}_{i}~{}\text{is not empty}),\\ 0;\;&\text{otherwise}.\end{cases}$ (25)
The first term in (24) is the attractive component of the potential function, which is active at all times. This term has a unique minimum at $e_{i}=0$ and steers the $i$th pursuer to the desired proximity around the target. The second term in (24) is the repulsive term enabling the $i$th pursuer to avoid collisions with its nearest colliding pursuer. It is important to note that the latter term is conditional and is only activated based on the existence of the nearest colliding pursuer for $P_{i}$, as per the condition given in (25). The first term in (24) is a quadratic term dependent on $e_{i}$, whereas the second term (which introduces a repulsive behavior) is a Gaussian function that attains maximum values at $e_{i}=e_{j}$ and decreases as $d_{ij}=e_{i}-e_{j}$ increases, with this term losing its dominance when $d_{ij}\geq\Delta_{i}$. If two agents are closer to each other, it is reasonable to assume $d_{ij}<e_{i}$. Therefore, the first term in (24) will dominate the second term for any selected value of $\lambda_{i}$, $\eta_{i}$ and $\Delta_{i}$ if $d_{ij}\gg 0$. However, for small values of $d_{ij}$, one can select the values of these parameters to ensure that the second term dominates in the region where $d_{ij}=e_{i}-e_{j}\in(0,\Delta_{i})$. In such a scenario, in the region near the line $e_{i}-e_{j}=0$, the repulsive term dominates the attractive component that will result in $P_{i}$ moving away from $P_{j}\in\mathcal{C}_{i}$. Consequently, it follows that there will be a shift in the minimum of the potential function at $e_{i}=0$ (for only attractive potential) to $e_{i}>e_{j}$ (for combined potential), which allows pursuers to avoid collisions with each other.
@@
$\frac{\partial U_{i}}{\partial e_{i}}$ $=\lambda_{i}e_{i}-\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j}),$ (26a)
$\frac{\partial U_{i}}{\partial e_{j}}$ $=\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j}).$ (26b)
To obtain the region of minimum potential, we equate (26a) to zero to obtain, that is, $\lambda_{i}e_{i}-\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j})=0.$ This above expression is highly nonlinear, and the analytical solution of $e_{i}$ is difficult to obtain in closed form. The trivial solution is $e_{i}=0$ and $e_{j}=0$. However, such a situation never arises because when $\delta_{i}=1$, $e_{i}>e_{j}$. Further, from the above equation, we can infer that the solution to $\dfrac{\partial U_{i}}{\partial e_{i}}=0$ and $\dfrac{\partial U_{i}}{\partial e_{j}}=0$, can be given by $e_{i}-e_{j}=\epsilon_{i}$, where $\epsilon_{i}$ denotes a positive value. To verify this, we plot the potential function and the negative gradients for the chosen (typical) values of $\lambda_{i}=0.2,\eta_{i}=100000$ and $\Delta_{i}=500$, as depicted in Figure 9. In Figure 9, $e_{i}-e_{j}>0$ is the relevant area whenever $\delta_{i}=1$. The combined potential is depicted in Figure 8(a), where it is observed that the minimum potential region (depicted by cyan colored region) is situated at $e_{i}-e_{j}=\epsilon_{i}$. Further, in Figure 8(b), the negative gradient lines converge to the line $e_{i}-e_{j}=\epsilon_{i}$. This depicts the existence of a minimum potential point at $e_{i}=e_{j}+\epsilon_{i}$, whenever the repulsive component of the potential is activated.
@@
It is crucial to choose the values of the control parameters $\lambda_{i}$, $\eta_{i}$, and $\Delta_{i}$ judiciously to guarantee that $\epsilon_{i}>0$. This ensures that the minimum of the combined potential function in (24) occurs at $e_{i}>e_{j}$, leading to $\dot{e}_{i}>0$ for all $e_{i}-e_{j}\in(0,\epsilon_{i})$ and $\dot{e}_{i}<0$ for all $e_{i}-e_{j}\in(\epsilon_{i},\infty)$. Based on (26a), the condition, $\dfrac{\delta_{i}}{\Delta_{i}^{2}\lambda_{i}}>\sup_{t\geq 0}\frac{e_{i}}{e_{i}-e_{j}}e^{\left(\frac{\left(e_{i}-e_{J}\right)^{2}}{2\Delta_{i}^{2}}\right)}$, will help select the control parameters to ensure $\epsilon_{i}>0$.
@@ III-C Development of the Target Enclosing Guidance Law
Differentiating $\dot{e}_{i}$ with respect to time and using (1b), we obtain the dynamics of range error rate as,
$\ddot{e}_{i}$ $=\ddot{r}_{iT}=\dfrac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+a_{i}\sin\sigma_{i},$ (27)
thereby completing the proof. ∎
@@
on which the $i$th pursuer’s range error rate $\dot{e}_{i}$ can be driven to the negative gradient $-\nabla_{e_{i}}\mathcal{U}_{ij}$. Note from the above equation that when $\delta_{i}=0$ $S_{i}=\dot{e}_{i}+\lambda_{i}e_{i}$, which is analogous to a second-order sliding manifold with asymptotic error convergence. Further, the last term in (28) represents the collision avoidance term that becomes active when $\delta_{i}=1$. We now propose the lateral acceleration (the only control input) for the $i$th pursuer as,
@@
$\textup{where},\,K_{i}>\frac{\eta_{i}v}{\Delta_{i}^{2}}\left(1-\frac{r_{s}^{2}}{\Delta_{i}^{2}}\right)$ (30)
provides a sufficient condition on the controller gain of the $i$th pursuer. The first term in the above equation drives the pursuers to their respective sliding manifold. The second term represents the centripetal acceleration about the target, while the third term represents the component of its velocity along its LOS with respect to the target. The fourth term in (III-C) enforces a collision avoidance behavior.
@@
Consider the equations of relative motion between a single target and multiple pursuers (II-A), and the $i$th pursuer-target range error $e_{i}$. If the $i$th pursuer applies the lateral acceleration given in (III-C), then all the pursuers will self-organize on circle of radius $r_{d}$ around target without colliding with each other.
@@
Consider a Lyapunov Function Candidate $V=\sum_{i\in\mathcal{L}}S_{i}^{2}.$ Differentiating this Lyapunov function with respect to time and using (28), we obtain the dynamics of the sliding manifold as,
$\dot{V}$ $=\sum_{i\in\mathcal{L}}S_{i}\dot{S}_{i}=\sum_{i\in\mathcal{L}}S_{i}\left(\ddot{e}_{i}+\lambda_{i}\dot{e}_{i}+\frac{\partial^{2}U_{i}}{\partial^{2}{e}_{i}}\dot{e}_{i}+\frac{\partial^{2}U_{i}}{\partial^{2}{e}_{j}}\dot{e}_{j}\right)$ (31)
Substituting for $\ddot{e}_{i},\,\partial^{2}\mathcal{U}_{i}/\partial^{2}e_{i}$ and $\partial^{2}\mathcal{U}_{i}/\partial^{2}e_{j}$ using (26a), (26b), and (27) in the above expression we obtain,
$\dot{V}$ $=\sum_{i\in\mathcal{L}}S_{i}\Biggr{[}\frac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+a_{i}\sin\sigma_{i}+\lambda\dot{e}_{i}$
$+\frac{\eta_{i}\delta_{i}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}}{\Delta_{i}^{2}}\left(\frac{({e}_{i}-{e}_{j})^{2}}{\Delta_{i}^{2}}-1\right)\left(\dot{e}_{i}-\dot{e}_{j}\right)\Biggr{]}.$ (32)
Substituting the proposed lateral acceleration for the $i$th pursuer given in (III-C), reduces the above expression to
@@
since $|\dot{e}_{j}|=v$ and $|e_{i}-e_{j}|<r_{s}$ (as the switching term can be activated only when the colliding vehicle is within the sensing range). The above equation can be further simplified as
$\dot{V}\leq\sum_{j\in\mathcal{L}}-|S_{i}|\Bigg{[}K_{i}-\frac{\eta_{i}v}{\Delta_{i}^{2}}\left(1-\frac{r_{s}^{2}}{\Delta_{i}^{2}}\right)\Bigg{]},$ (33)
@@
which presents the sufficient condition as given in (30) to ensure, $\dot{V}<0,\,\forall\,S_{i}\in\mathbb{R}\setminus\{0\}$. This implies that $S_{i}$’s converge to zero within finite time given by $T^{*}=|S_{i}(0)|/K_{i}$, irrespective whether $\delta_{i}=0\,\textup{or}\,1$. It is worth mentioning that $S_{i}$ can suddenly deviate from zero due to the activation/deactivation of the repulsive potential component. However, based on (33), even when $S_{i}$ shifts suddenly to a non-zero value, $S_{i}$ will converge to a zero within a finite time to quickly enforce the sliding mode. Once $S_{i}=0$, the pursuers’ error rate propagates according to
${\dot{e}}_{i}=-\lambda_{i}{e}_{i}+\frac{\eta_{i}\delta_{i}}{\Delta_{i}^{2}}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta_{i}^{2}}}({e}_{i}-{e}_{j})=0.$ (34)
Therefore, once the sliding mode is enforced, $P_{i}$ may exhibit two distinct behaviors in relation to other vehicles based on the switching term, discussed as Modes 1 and 2 below.
@@
Mode 1: When $\delta_{i}=0$, indicating the absence of any nearest colliding pursuer for $P_{i}$, then the dynamics described in (34) yields ${\dot{e}}_{i}=-\lambda_{i}{e}_{i}$. This leads to 1) ${\dot{e}}_{i}\leq 0$ when $e_{i}\geq 0$ and 2) ${\dot{e}}_{i}>0$ when $e_{i}<0$. The solution of the range error can obtained as $\;{e}_{i}(t)={e}_{i}(0)e^{-\lambda_{i}t}$. Consequently, the pursuer converges to the desired proximity, that is, $r_{iT}\rightarrow r_{d}$ if $\lambda_{i}$ is selected as a positive value.
Mode II: If ${\delta}_{i}=1$, this pertains to the scenario when there exists a nearest colliding neighbor $P_{j}$ for the pursuer $P_{i}$ such that $d_{ij}>0$ and $\psi_{ij}<0$ at the start of the engagement. In this mode, (34) yields
${\dot{e}}_{i}=-\lambda_{i}{e}_{i}+\frac{\eta_{i}}{\Delta^{2}}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}({e}_{i}-{e}_{j})=0,$ (35)
@@
as $\delta_{i}=1$. It can be inferred that the solution of $e_{i}(t)$ when $\delta_{i}=1$ may not be feasible to obtain analytically in each case. However, we can select the values of controller parameters $\lambda_{i}$, $\eta_{i}$ and $\Delta_{i}$ to ensure $\dot{e_{i}}>0$ when $e_{i}-e_{j}<\epsilon_{i}$ and $\dot{e_{i}}<0$ when $e_{i}-e_{j}>\epsilon_{i}$, where $\epsilon_{i}>0$ denotes a positive value. This is similar to the behavior of the pursuer under the proposed potential function as discussed earlier in Section III-B. This implies $P_{i}$ will converge to a loiter circle of radius greater than that of $P_{j}$, that is, $e_{i}\to e_{j}+\epsilon_{i}\,\,\textup{or}\,\,r_{iT}\to r_{jT}+\epsilon_{i}.$ It readily follows from the above relation that $d_{ij}>0$, which represents the safety condition as in (8). The above condition ensures $d_{ij}>0,\,\,\forall\,t\geq 0$, while $\psi_{ij}<0$ and $\dot{\psi}_{ij}>0$ in Region IV of the $i$th pursuer, as obtained in (III-A) of Lemma 1. This will result in $\psi_{ij}$ starting from a negative value to become zero. At this moment, $P_{j}$ will no longer be the nearest colliding neighbor for $P_{i}$. Further, if no other vehicle becomes the new nearest colliding neighbor, $P_{i}$ will undergo the behavior in Mode I to reach the desired proximity, settling behind $P_{j}$ and exhibiting the self-organizing behavior. Otherwise, $P_{i}$ will undergo the behavior in Mode II to avoid collision with the new nearest colliding neighbor and settle behind it. Consequently, every pursuer starting away from the target will alternate between Mode I and Mode II behaviors, resulting in every pursuer to converge to circle of desired proximity without colliding with other pursuers and organising autonomously on enclosing shape. This concludes the proof. ∎
@@
Ensuring the expression of $K_{i}$ as in (30) is greater than zero provides an additional condition on the selection of the control parameters, that is, $\Delta_{i}>r_{s}$. This simply implies that the region of the dominance of the repulsive term denoted by $\Delta_{i}$ is greater than the $i$th pursuer’s sensing radius, resulting in the repulsive field dominating within the sensing radius. This will enable the pursuers to move out of the sensing range of their respective nearest colliding pursuers to avoid colliding with them, such that at steady state, none of the pursuers are situated in each other’s sensing radius.

### v2

@@ III-B Design of Potential Function
As indicated by the behavior of pursuers observed in Section III-A, each pursuer must exhibit two distinct behaviors. When there is no nearby colliding pursuer, the $i$th pursuer’s objective is to reach the desired proximity from the target to enclose it. However, if there is a nearest colliding pursuer, the $i$th pursuer needs to maneuver along a loiter circle with a radius larger than the one corresponding to the nearest colliding pursuer to prevent any potential collisions. These factors enable us to break down the complex multiagent target enclosing problem into more manageable two-body (one pursuer and the target) or three-body (two pursuers and the target) scenarios at any given point in time. Consequently, in this subsection, we present the design of a potential function that will serve as the foundation for developing the target enclosing guidance law.
Let’s define the range error for the $i$th pursuer as, ${e}_{i}=r_{iT}-r_{d},$ where $r_{d}$ is the desired fixed proximity from the target such that $\dot{r}_{d}=0$. This error variable plays a crucial role in the design of the potential function to ensure inter-pursuer safety. To analyze the error dynamics, we differentiate $e_{i}$ with respect to time and use (1a) to obtain the range error rate for the $i$th pursuer as, $\dot{e}_{i}=\dot{r}_{iT}=-v\cos\sigma_{i}$, since $\dot{r}_{d}=0$. From the above equation, it is evident that $\dot{e}_{i}<0,$ if $\sigma_{i}\in[0,\pi/2)$, which indicates the $i$th pursuer’s motion toward the target, whereas $\dot{e}_{i}>0,$ represents the opposite case because $\sigma_{i}\in(\pi/2,\pi]$. The $i$th pursuer moves on a circle of fixed radius from the target when $\dot{e}_{i}=0$ when $\sigma_{i}=\pi/2$. Therefore, nullifying the range error $e_{i}$ will ensure that the $i$th pursuer converges to the desired proximity around the target, but additional safety considerations are needed to prevent the inter-pursuer collision.
@@
$\mathcal{U}_{i}(e_{i},e_{j})=\lambda_{i}\frac{e_{i}^{2}}{2}+\eta_{i}\delta_{i}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}$ (13)
where $\lambda_{i}$, $\eta_{i}$, and $\Delta_{i}$ are positive scaling constants in the potential function. The subscript $j$ corresponds to $P_{j}\in\mathcal{C}_{i}$, which is the nearest colliding pursuer corresponding to the $i$th pursuer. The variable $\delta_{i}$ is a conditional one such that $\delta_{i}=\begin{cases}1;\,&\text{if}\,\mathcal{C}_{i}\neq\emptyset~{}(\mathcal{C}_{i}~{}\text{is not empty}),\\ 0;\;&\text{otherwise}.\end{cases}$. The first term in (13) is the attractive component of the potential function, which is active at all times. This term has a unique minimum at $e_{i}=0$ and steers the $i$th pursuer to the desired proximity around the target. The second term in (13) is the repulsive term enabling the $i$th pursuer to avoid collisions with its nearest colliding pursuer. It is important to note that the latter term is conditional and is only activated based on the existence of the nearest colliding pursuer for $P_{i}$. The first term in (13) is a quadratic term dependent on $e_{i}$, whereas the second term (which introduces a repulsive behavior) is a Gaussian function that attains maximum values at $e_{i}=e_{j}$ and decreases as $d_{ij}=e_{i}-e_{j}$ increases, with this term losing its dominance when $d_{ij}\geq\Delta_{i}$. The Gaussian function in (13) is chosen as the potential function for collision avoidance due to its localized influence, characterized by an exponential decrease in value with distance from the mean. This function’s shape can be precisely controlled through its parameters, such as the mean and standard deviation, offering significant adaptability. In comparison, alternative repulsive potential functions, such as inverse proportional or logarithmic functions, exhibit non-smooth behavior, reduced adaptability, and a rapid decrease to zero as the distance from the mean increases. These attributes may render them less suitable for effective collision avoidance in a dynamically changing environment. If two agents are closer to each other, it is reasonable to assume $d_{ij}<e_{i}$. Therefore, the first term in (13) will dominate the second term for any selected value of $\lambda_{i}$, $\eta_{i}$ and $\Delta_{i}$ if $d_{ij}\gg 0$. However, for small values of $d_{ij}$, one can select the values of these parameters to ensure that the second term dominates in the region where $d_{ij}=e_{i}-e_{j}\in(0,\Delta_{i})$. In such a scenario, in the region near the line $e_{i}-e_{j}=0$, the repulsive term dominates the attractive component that will result in $P_{i}$ moving away from $P_{j}\in\mathcal{C}_{i}$. Consequently, it follows that there will be a shift in the minimum of the potential function at $e_{i}=0$ (for only attractive potential) to $e_{i}>e_{j}$ (for combined potential), which allows pursuers to avoid collisions with each other.
@@
$\frac{\partial\mathcal{U}_{i}}{\partial e_{i}}$ $=\lambda_{i}e_{i}-\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j}),$ (14a)
$\frac{\partial{\mathcal{U}_{i}}}{\partial e_{j}}$ $=\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j}).$ (14b)
To obtain the region of minimum potential, we equate (14a) to zero to obtain, that is, $\lambda_{i}e_{i}-\frac{\eta_{i}\delta_{i}}{\Delta^{2}}e^{-\frac{(e_{i}-e_{j})^{2}}{2\Delta_{i}^{2}}}(e_{i}-e_{j})=0.$ This above expression is highly nonlinear, and the analytical solution of $e_{i}$ is difficult to obtain in closed form. The trivial solution is $e_{i}=0$ and $e_{j}=0$. However, such a situation never arises because when $\delta_{i}=1$, $e_{i}>e_{j}$. Further, from the above equation, we can infer that the solution to $\dfrac{\partial{\mathcal{U}_{i}}}{\partial e_{i}}=0$ and $\dfrac{\partial{\mathcal{U}_{i}}}{\partial e_{j}}=0$, can be given by $e_{i}-e_{j}=\epsilon_{i}$, where $\epsilon_{i}$ denotes a positive value. To verify this, we plot the potential function and the negative gradients for the chosen (typical) values of $\lambda_{i}=0.2,\eta_{i}=100000$ and $\Delta_{i}=500$, as shown in Figure 6. In Figure 6, $e_{i}-e_{j}>0$ is the relevant area whenever $\delta_{i}=1$. The combined potential is depicted in Figure 6(a), where it is observed that the minimum potential region (depicted by cyan colored region) is situated at $e_{i}-e_{j}=\epsilon_{i}$. Further, in Figure 6(b), the negative gradient lines converge to the line $e_{i}-e_{j}=\epsilon_{i}$. This depicts the existence of a minimum potential point at $e_{i}=e_{j}+\epsilon_{i}$, when the repulsive component of the potential is activated.
@@
It is crucial to choose the values of the control parameters $\lambda_{i}$, $\eta_{i}$, and $\Delta_{i}$ judiciously to guarantee that $\epsilon_{i}>0$. This ensures that the minimum of the combined potential function in (13) occurs at $e_{i}>e_{j}$, leading to $\dot{e}_{i}>0$ for all $e_{i}-e_{j}\in(0,\epsilon_{i})$ and $\dot{e}_{i}<0$ for all $e_{i}-e_{j}\in(\epsilon_{i},\infty)$. To ensure that the repulsive term in potential function gradient in (14a) dominates whenever activated, the condition $\dfrac{\delta_{i}}{\Delta_{i}^{2}\lambda_{i}}>\sup_{t\geq 0}\frac{e_{i}}{e_{i}-e_{j}}\exp\left\{\left(\frac{\left(e_{i}-e_{J}\right)^{2}}{2\Delta_{i}^{2}}\right)\right\}$ is obtained, which will help select the control parameters to ensure $\epsilon_{i}>0$.
@@ III-C Development of the Target Enclosing Guidance Law
Differentiating $\dot{e}_{i}$ with respect to time and using (1b), we obtain the dynamics of range error rate as, $\ddot{e}_{i}=\ddot{r}_{iT}=\dfrac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+a_{i}\sin\sigma_{i},$ thereby completing the proof. ∎
@@
on which the $i$th pursuer’s range error rate $\dot{e}_{i}$ can be driven to the negative gradient $-\nabla_{e_{i}}{\mathcal{U}_{i}}$. Note from the above equation that when $\delta_{i}=0$ $S_{i}=\dot{e}_{i}+\lambda_{i}e_{i}$, which is analogous to a second-order sliding manifold with asymptotic error convergence. Further, the last term in (15) represents the collision avoidance term that becomes active when $\delta_{i}=1$. We now propose the lateral acceleration (the only control input) for the $i$th pursuer as,
@@
where $K_{i}>\frac{\eta_{i}v}{\Delta_{i}^{2}}\left(1-\frac{{r_{s}}^{2}}{\Delta_{i}^{2}}\right)$ provides a sufficient condition on the controller gain of the $i$th pursuer. The first term in the above equation drives the pursuers to their respective sliding manifold. The second term represents the centripetal acceleration about the target, while the third term represents the component of its velocity along its LOS with respect to the target. The fourth term in (III-C) enforces a collision avoidance behavior.
@@
Consider the equations of relative motion between a single target and multiple pursuers (1), and the $i$th pursuer-target range error $e_{i}$. If the $i$th pursuer applies the lateral acceleration given in (III-C), then all the pursuers will self-organize on a circle of radius $r_{d}$ around target without colliding with each other.
@@
Consider a Lyapunov Function Candidate $V=\sum_{i\in\mathcal{L}}{\frac{1}{2}}S_{i}^{2}.$ Differentiating this Lyapunov function with respect to time and using (15), we obtain the dynamics of the sliding manifold as, $\dot{V}=\sum_{i\in\mathcal{L}}S_{i}\dot{S}_{i}=\sum_{i\in\mathcal{L}}S_{i}\left(\ddot{e}_{i}+\lambda_{i}\dot{e}_{i}+\frac{\partial^{2}\mathcal{U}_{i}}{\partial^{2}{e}_{i}}\dot{e}_{i}+\frac{\partial^{2}\mathcal{U}_{i}}{\partial^{2}{e}_{j}}\dot{e}_{j}\right).$ Substituting for $\ddot{e}_{i},\,\partial^{2}\mathcal{U}_{i}/\partial^{2}e_{i}$ and $\partial^{2}\mathcal{U}_{i}/\partial^{2}e_{j}$ in the above expression we obtain $\dot{V}=\sum_{i\in\mathcal{L}}S_{i}\Biggr{[}\frac{v^{2}\sin^{2}{\sigma_{i}}}{r_{iT}}+a_{i}\sin\sigma_{i}+\lambda\dot{e}_{i}+\frac{\eta_{i}\delta_{i}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}}{\Delta_{i}^{2}}\left(\frac{({e}_{i}-{e}_{j})^{2}}{\Delta_{i}^{2}}-1\right)\left(\dot{e}_{i}-\dot{e}_{j}\right)\Biggr{]}.$ Substituting the proposed lateral acceleration for the $i$th pursuer given in (III-C), reduces the above expression to
@@
since $|\dot{e}_{j}|=v$ and $|e_{i}-e_{j}|<r_{s}$ (as the switching term can be activated only when the colliding vehicle is within sensing range). The above equation can be further simplified as, $\dot{V}\leq\sum_{j\in\mathcal{L}}-|S_{i}|\Bigg{[}K_{i}-\frac{\eta_{i}v}{\Delta_{i}^{2}}\left(1-\frac{{r_{s}}^{2}}{\Delta_{i}^{2}}\right)\Bigg{]},$ which presents the sufficient condition as given after (III-C) to ensure, $\dot{V}<0,\,\forall\,S_{i}\in\mathbb{R}\setminus\{0\}$. This implies that $S_{i}$’s converge to zero within finite time given by $T^{*}=|S_{i}(0)|/K_{i}$, irrespective whether $\delta_{i}=0\,\textup{or}\,1$. It is worth mentioning that $S_{i}$ can suddenly deviate from zero due to the activation/deactivation of the repulsive potential component. However, based on our control law, even when $S_{i}$ shifts suddenly to a non-zero value, $S_{i}$ will converge to a zero within a finite time to quickly enforce the sliding mode. Once $S_{i}=0$, the pursuers’ error rate propagates according to ${\dot{e}}_{i}=-\lambda_{i}{e}_{i}+\frac{\eta_{i}\delta_{i}}{\Delta_{i}^{2}}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta_{i}^{2}}}({e}_{i}-{e}_{j})=0.$ Therefore, once the sliding mode is enforced, $P_{i}$ may exhibit two distinct behaviors in relation to other vehicles based on the switching term, discussed as Modes 1 and 2 below.
@@
Mode 1: When $\delta_{i}=0$, indicating the absence of any nearest colliding pursuer for $P_{i}$, then the error dynamics is described as ${\dot{e}}_{i}=-\lambda_{i}{e}_{i}$ (since $\delta_{i}=0$). This leads to 1) ${\dot{e}}_{i}\leq 0$ when $e_{i}\geq 0$ and 2) ${\dot{e}}_{i}>0$ when $e_{i}<0$. The solution of the range error can obtained as $\;{e}_{i}(t)={e}_{i}(0)e^{-\lambda_{i}t}$. Consequently, the pursuer converges to the desired proximity, that is, $r_{iT}\rightarrow r_{d}$ if $\lambda_{i}$ is selected as a positive value.
@@
Mode II: If ${\delta}_{i}=1$, this pertains to the scenario when there exists a nearest colliding neighbor $P_{j}$ for the pursuer $P_{i}$ such that $d_{ij}>0$ and $\psi_{ij}<0$ at the start of the engagement. the error dynamics is described as ${\dot{e}}_{i}=-\lambda_{i}{e}_{i}+\frac{\eta_{i}}{\Delta^{2}}e^{-\frac{({e}_{i}-{e}_{j})^{2}}{2\Delta^{2}}}({e}_{i}-{e}_{j})=0,$ as $\delta_{i}=1$. It can be inferred that the solution of $e_{i}(t)$ when $\delta_{i}=1$ may not be feasible to obtain analytically in each case. However, we can select the values of controller parameters $\lambda_{i}$, $\eta_{i}$ and $\Delta_{i}$ to ensure $\dot{e_{i}}>0$ when $e_{i}-e_{j}<\epsilon_{i}$ and $\dot{e_{i}}<0$ when $e_{i}-e_{j}>\epsilon_{i}$, where $\epsilon_{i}>0$ denotes a positive value. This is similar to the behavior of the pursuer under the proposed potential function as discussed earlier in Section III-B. This implies $P_{i}$ will converge to a loiter circle of radius greater than that of $P_{j}$, that is, $e_{i}\to e_{j}+\epsilon_{i}\,\,\textup{or}\,\,r_{iT}\to r_{jT}+\epsilon_{i}.$ It readily follows from the above relation that $d_{ij}>0$, which represents the safety condition as in (8). The above condition ensures $d_{ij}>0,\,\,\forall\,t\geq 0$, while $\psi_{ij}<0$ and $\dot{\psi}_{ij}>0$ in Region IV of the $i$th pursuer, as obtained in (III-A) of Lemma 1. This will result in $\psi_{ij}$ starting from a negative value to become zero. At this moment, $P_{j}$ will no longer be the nearest colliding neighbor for $P_{i}$. Further, if no other vehicle becomes the new nearest colliding neighbor, $P_{i}$ will undergo the behavior in Mode I to reach the desired proximity, settling behind $P_{j}$ and exhibiting the self-organizing behavior. Otherwise, $P_{i}$ will undergo the behavior in Mode II to avoid collision with the new nearest colliding neighbor and settle behind it. Consequently, every pursuer starting away from the target will alternate between Mode I and Mode II behaviors, resulting in every pursuer to converge to circle of desired proximity without colliding with other pursuers and organising autonomously on enclosing shape. This concludes the proof. ∎
@@
Ensuring the expression of $K_{i}$ as after (III-C) is greater than zero provides an additional condition on the selection of the control parameters, that is, $\Delta_{i}>r_{s}$. This simply implies that the region of the dominance of the repulsive term denoted by $\Delta_{i}$ is greater than the $i$th pursuer’s sensing radius, resulting in the repulsive field dominating within the sensing radius. Additionally, the selected values of parameters $\Delta_{i},\delta_{i}$ and $\lambda_{i}$ directly impacts the minimum allowable safe distance $\epsilon_{i}$ (as discussed in Remark 6) among the pursuers. A larger value of $\epsilon_{i}$ results in a lesser number of pursuers safely converging to the desired proximity and vice-versa at the steady state.